# Safe and Economical UAV Trajectory Planning in Low-Altitude Airspace: A Hybrid DRL-LLM Algorithm with Compliance Awareness

### v2

@@ 4.3 Problem Characteristics
This optimization problem involves continuous control, nonlinear constraints, dynamic environmental factors (such as OUs), as well as regulatory and energy limitations, making it structurally complex and difficult. Traditional optimization algorithms often perform inefficiently and lack adaptability when dealing with such high-dimensional, non-convex, dynamic, and partially observable problems. In contrast, DRL offers adaptive learning capabilities, making it well-suited for continuous control and long-term optimization. LLMs, on the other hand, can provide rule-compliant and reasoning decisions in critical scenarios by leveraging rich prior knowledge. The integration of DRL and LLM leverages efficiency, generalization, and safety, making it more suitable than traditional methods to solve complex UAV trajectory planning problems.
@@ 5.1 POMDP Modeling
Observation Space $\mathcal{O}$: $o_{t}\in\mathcal{O}$ is the state that the DCU can observe at time $t$, where $o_{t}=(o_{GE,t},o_{Z,t},o_{u^{\prime},t},o_{u,t}).$
  • $o_{GE,t}$ is the state about all GEs and $o_{GE,t}=\{L_{GE,i},DV_{GE_{i},t},TP_{GE,i}\}_{i=1}^{N_{GE}}$.
  • $o_{Z,t}$ is the state about all zones and $o_{Z,t}=\{\{\mathbb{R}_{NFZ,l}\}_{l=1}^{N_{NFZ}},\{\mathbb{R}_{RZ,m}\}_{m=1}^{N_{RZ}},\{\mathbb{R}_{BZ,n}\}_{n=1}^{N_{BZ}}\}$.
  • $o_{u^{\prime},t}$ is the state about all OUs and $o_{u^{\prime},t}=\{L_{u^{\prime}_{j},t},v_{u^{\prime}_{j},t}\}_{j=1}^{N_{U^{\prime}}}$.
  • $o_{u,t}$ is the state about DCU itself and $o_{u,t}=\{L_{u,t},v_{u,t},E(t),TD_{u,GE_{i}}(t),\mathbb{T}\mathbb{A},\mathbb{L}\mathbb{A}\}$.
Action Space $\mathcal{A}$: $a_{t}\in\mathcal{A}$ is the sampled action at time $t$ and $a_{t}=(v_{x}(t),v_{y}(t),0),$ which means the DCU will move at the velocity $(v_{x}(t),v_{y}(t),0)$ at time $t$.
@@
${r_{1t}}=\left\{\begin{array}[]{ll}{\sigma_{1}},&\text{if }TD_{u,GE_{i}}(t+1)\geq 0,\\ 0,&\text{otherwise}.\end{array}\right$.
@@
${r_{2t}}=\left\{\begin{array}[]{ll}0,&\text{if }d_{t,u^{\prime}}\geq d_{safe,u^{\prime}},\\ -\sigma_{2}\times\frac{d_{safe,u^{\prime}}-d_{t,u^{\prime}}}{d_{safe,u^{\prime}}-d_{\min,u^{\prime}}},&\text{if }d_{\min,u^{\prime}}<d_{t}<d_{safe,u^{\prime}},\\ -\sigma_{2},&\text{if }d_{t,u^{\prime}}\leq d_{\min,u^{\prime}},\end{array}\right$.
where ${d_{t,u^{\prime}}}=\arg\min\{||L_{u,t}-L_{u^{\prime}_{j},t}||\}_{j=1}^{N_{U^{\prime}}}$, $d_{\min,u^{\prime}}$} is the minimum safe distance between the DCU and OUs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,u^{\prime}}$} is the safe distance threshold between the DCU and OUs.
@@
where ${d_{t,BZ}}=\arg\min\{||L_{u,t}-\mathbb{R}_{BZ,n}||\}_{n=1}^{N_{BZ}}$, $d_{\min,BZ}$} is the minimum safe distance between the DCU and tall buildings, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,BZ}$} is the safe distance threshold between the DCU and tall buildings.
@@
where $d_{t,NFZ}=\arg\min\left\{\|L_{u,t}-\mathbb{R}_{NFZ,l}\|\right\}_{l=1}^{N_{NFZ}}$, $d_{\min,NFZ}$ is the minimum safe distance between the DCU and NFZs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{safe,NFZ}$ is the safe distance threshold between the DCU and NFZs.
@@
where $d_{t,\mathbb{A}}=\|L_{u,t}-\mathbb{A}\|$ is the distance between the DCU and the boundary of $\mathbb{A}$ at time $t$, $d_{\min,\mathbb{A}}$ is the minimum safe distance between the DCU and the boundary of $\mathbb{A}$, which is used for the DCU to adjust the speed to avoid flying out, and $d_{safe,\mathbb{A}}$ is the safe distance threshold between the DCU and the boundary of $\mathbb{A}$.
@@
where $e_{\min}(t)=e_{\text{fly}}(t)\cdot d_{t,\mathbb{LA}}/v_{\max}$.
$r_{9t}$ is used to encourage the DCU can complete the data collection task quickly to reduce energy consumption and is as follows:
@@ 5.2 SAC Algorithm
SAC is a state-of-the-art off-policy DRL algorithm based on the maximum entropy framework [39]. Its off-policy nature allows flexible integration of external guidance from the LLM without compromising training stability or policy optimization efficiency. Additionally, SAC is well-suited for our problem, which involves a continuous action space. Although we adopt SAC in this study, we believe that other off-policy DRL algorithms supporting continuous action spaces may also achieve comparable performance, which is detailed in Section 6.
The maximum entropy DRL incorporates the policy’s entropy into the objective function [40], modifying the learning goal as follows:
$\pi^{*}=\arg\max_{\pi}\sum_{t=0}^{\infty}\mathbb{E}_{\mathcal{P}\pi}\left[\delta^{t}\left(r_{t}+\mu P(\pi(\cdot\mid o_{t}))\right)\right],$ (20)
where $\delta$ is the discounted factor and $\mu$ is the temperature parameter. The policy entropy $P(x)$ is defined as follows:
$P(x)=-\sum\nolimits_{x}p(x)\log p(x).$ (21)
Given the soft value-state function and the soft value function as follows:
$Q_{\text{soft}}(o_{t},a_{t})=r_{t}+\mathbb{E}_{\mathcal{P}\pi}\left[\sum\nolimits_{i=1}^{\infty}\delta^{i}\left(r_{t+i}+\mu P(\cdot\mid\pi)\right)\right],$ (22)
$V_{\text{soft}}(o_{t})=\mu\log\int_{\mathcal{A}}\exp\left(\frac{Q_{\text{soft}}(o_{t},a_{t+1})}{\mu}\right)da_{t+1}.$ (23)
They can be iterated in the following ways:
$Q_{\text{soft}}(o_{t},a_{t})=r_{t}+\delta\mathbb{E}_{\mathcal{P}}\left[V_{\text{soft}}(o_{t+1})\right],$ (24)
$V_{\text{soft}}(o_{t})=\mathbb{E}_{\pi}\left[Q_{\text{soft}}(o_{t},a_{t})-\mu\log\pi(a_{t}\mid o_{t})\right].$ (25)
Therefore, the policy $\pi$ is as follows:
$\pi(o_{t},a_{t})=\exp\left(\frac{Q_{\text{soft}}(o_{t},a_{t})}{\delta}-V_{\text{soft}}(o_{t})\right).$ (26)
In SAC, both $\pi$ and $Q_{\text{soft}}$ are parameterized using neural networks. This enables SAC to handle continuous state and action spaces, making it well-suited for solving complex problems and enhancing controller performance in scenarios, such as trajectory planning. The neural networks associated with $Q_{\text{soft}}$ are updated at each step by minimizing the following loss function:
$J_{Q}(\varphi_{i})=\mathbb{E}_{(o_{t},a_{t},r_{t},o_{t+1})\sim D}\left[\left(Q_{i}^{\theta}(o_{t},a_{t})-y(r_{t},o_{t+1})\right)^{2}\right],$ (27)
where $y(r_{t},o_{t+1})$ is as follows:
$y(r_{t},o_{t+1})=$ (28)
$r_{t}+\delta\left(\min_{j=1,2}Q_{i}^{\theta_{\text{tar}}}(o_{t+1},\tilde{a}_{t+1})-\mu\log\pi^{\phi}(\tilde{a}_{t+1}\mid o_{t+1})\right).$
For $\pi$, with the sampled action, it is optimized as follows:
$J_{\pi}(\phi)=$ $\mathbb{E}_{o_{t}\sim D,\varsigma\sim N}\big{[}\mu\log\pi^{\phi}(\tilde{a}_{t}^{\phi}(o_{t},\varsigma)\mid o_{t})$- (29)
$\min_{j=1,2}Q_{i}^{\theta_{\text{tar}}}(o_{t},\tilde{a}_{t}^{\phi}(o_{t},\varsigma))\big{]},$
where $\tilde{a}_{t}^{\phi}(o_{t},\varsigma)$ is as follows:
$\tilde{a}_{t}^{\phi}(o_{t},\varsigma)=\tanh\left(\beta^{\phi}(o_{t})+\alpha^{\phi}(o_{t})\odot\varsigma\right),\quad\varsigma\sim\mathcal{N}(0,1),$ (30)
where $\beta$ and $\alpha$ represent the mean and standard deviation of the Gaussian distribution, respectively; $\varsigma$ denotes the noise; and $\mathcal{N}(0,1)$ indicates the standard normal distribution.
@@ 5.3.1 Main Idea
Based on the POMDP formulation in Section 5.1 and the continuous nature of DCU’s velocity, we adopt SAC algorithm to optimize the DCU’s flight trajectory during the data collection task. However, in the context of low-altitude economy scenarios, where strict regulatory compliance and robust obstacle-avoidance capability are critical, the use of DRL alone may be insufficient due to its reliance on trial-and-error learning and lack of prior domain knowledge. To address this issue, we propose an enhanced decision-making algorithm by integrating an LLM into the action selection module, as illustrated in Fig.  2. Specifically, when the DCU detects the presence of nearby obstacles (e.g., BZs, NFZs, RZs, or OUs), a decision condition is triggered that activates the LLM-based action selection pathway. In this mode, a Prompt Generator compiles semantic information, including environmental context, DCU status, and obstacle details, into a structured prompt, which is then input to an LLM. The LLM responds with a recommended action, leveraging its rich prior knowledge to ensure safety and regulatory compliance. An Extractor then parses the LLM’s response into actionable control parameters for the DCU. Conversely, when no immediate obstacle is detected, the system defaults to the standard SAC policy network for action selection. This hybrid structure enables the agent to benefit from both data-driven learning and knowledge-driven reasoning: the SAC component provides efficient trajectory planning in general scenarios, while the LLM supplements decision-making in critical or unfamiliar environments requiring sophisticated judgment. By combining the generalization capability of DRL with the reasoning capability and prior knowledge of an LLM, the proposed algorithm achieves superior performance in complex real-world scenarios, ensuring not only energy efficiency but also operational safety and compliance.
@@ 5.3.2 Algorithm Detail
Our proposed algorithm, which is detailed in Algorithm 1, is designed to enable the DCU to effectively perform data collection by learning an optimal velocity control policy that guides its trajectory to maximize data acquisition and avoid collision. The main process of our algorithm is shown in Fig. 2.

### v3

@@ 4.3 Problem Characteristics
This optimization problem involves continuous control, nonlinear constraints, dynamic environmental factors (such as OUs), as well as regulatory and energy limitations, making it structurally complex and difficult. Traditional optimization algorithms often perform inefficiently and lack adaptability when dealing with such high-dimensional, dynamic, and partially observable problems. In contrast, DRL offers adaptive learning capabilities, making it well-suited for continuous control and long-term optimization. LLMs, on the other hand, can provide rule-compliant and reasoning decisions in critical scenarios by leveraging rich prior knowledge. The integration of DRL and LLM leverages efficiency, generalization, and safety, making it more suitable than traditional methods to solve complex UAV trajectory planning problems.
@@ 5.1 POMDP Modeling
Observation Space $\mathcal{O}$: $o_{t}\in\mathcal{O}$ is the state that the DCU can observe at time $t$, where $o_{t}=(o_{\mathrm{GE},t},o_{Z,t},o_{\mathrm{OU},t},o_{\mathrm{DCU},t}).$
  • $o_{\mathrm{GE},t}$ is the state about all GEs and $o_{\mathrm{GE},t}=\{L_{{\mathrm{GE}}_{i}},DV_{\mathrm{GE}_{i},t},TP_{{\mathrm{GE}}_{i}}\}_{i=1}^{N_{\mathrm{GE}}}$.
  • $o_{Z,t}$ is the state about all zones and $o_{Z,t}=\{\{\mathbb{R}_{\mathrm{NFZ},l}\}_{l=1}^{N_{\mathrm{NFZ}}},\{\mathbb{R}_{\mathrm{RZ},m}\}_{m=1}^{N_{\mathrm{RZ}}},\{\mathbb{R}_{\mathrm{BZ},n}\}_{n=1}^{N_{\mathrm{BZ}}}\}$.
  • $o_{\mathrm{OU},t}$ represents the state information of OUs located within the perception radius of the DCU during time slot $t$. It is defined as $o_{\mathrm{OU},t}=\{(L_{\mathrm{OU}_{j},t},v_{\mathrm{OU}_{j},t})\}_{j=1}^{N_{\mathrm{OU}}}$, where $L_{\mathrm{OU}_{j},t}$ and $v_{\mathrm{OU}_{j},t}$ denote the position and velocity of $\mathrm{OU}_{j}$ at the beginning of time slot $t$. The number of OUs within the perception radius of the DCU may vary across different time slots, resulting in changes in the dimension of $o_{\mathrm{OU},t}$. To maintain a consistent input size, the zero-padding method is employed to standardize the dimension of $o_{\mathrm{OU},t}$.
  • $o_{\mathrm{DCU},t}$ is the state about DCU itself and $o_{\mathrm{DCU},t}=\{L_{\mathrm{DCU},t},v_{\mathrm{DCU},t},E(t),TD_{\mathrm{GE}_{i}}(t),\mathbb{T}\mathbb{A},\mathbb{L}\mathbb{A}\}$.
Action Space $\mathcal{A}$: $a_{t}\in\mathcal{A}$ is the sampled action at time $t$ and $a_{t}=(v_{x}(t),v_{y}(t)),$ which means the DCU will move at the velocity $(v_{x}(t),v_{y}(t))$ at time $t$.
@@
${r_{1t}}=\left\{\begin{array}[]{ll}{\sigma_{1}},&\text{if }TD_{\mathrm{GE}_{i}}(t+1)\geq 0,\\ 0,&\text{otherwise}.\end{array}\right$. (11)
@@
$r_{2t}=\left\{\begin{array}[]{ll}0,&\text{if }d_{t,\mathrm{OU}}\geq d_{\mathrm{safe},\mathrm{OU}},\\ -\sigma_{2}\displaystyle\frac{d_{\mathrm{safe},\mathrm{OU}}-d_{t,\mathrm{OU}}}{d_{\mathrm{safe},\mathrm{OU}}-d_{\min,\mathrm{OU}}},&\begin{aligned} &\text{if }d_{\min,\mathrm{OU}}<d_{t,\mathrm{OU}}.\\[-3.44444pt] &\text{and }d_{t,\mathrm{OU}}<d_{\mathrm{safe},\mathrm{OU}},\end{aligned}\\[3.44444pt] -\sigma_{2},&\text{if }d_{t,\mathrm{OU}}\leq d_{\min,\mathrm{OU}},\end{array}\right$. (12)
where ${d_{t,\mathrm{OU}}}=\arg\min\{||L_{\mathrm{DCU},t}-L_{\mathrm{OU}_{j},t}||\}_{j=1}^{N_{\mathrm{OU}}}$, $d_{\min,\mathrm{OU}}$} is the minimum safe distance between the DCU and OUs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{\mathrm{safe},\mathrm{OU}}$} is the safe distance threshold between the DCU and OUs.
@@
where ${d_{t,\mathrm{BZ}}}=\arg\min\{||L_{\mathrm{DCU},t}-\mathbb{R}_{\mathrm{BZ},n}||\}_{n=1}^{N_{\mathrm{BZ}}}$, $d_{\min,\mathrm{BZ}}$} is the minimum safe distance between the DCU and tall buildings, which is used for the DCU to adjust the speed to avoid collisions, and $d_{\mathrm{safe},\mathrm{BZ}}$} is the safe distance threshold between the DCU and tall buildings.
@@
where $d_{t,\mathrm{NFZ}}=\arg\min\left\{\|L_{\mathrm{DCU},t}-\mathbb{R}_{\mathrm{NFZ},l}\|\right\}_{l=1}^{N_{\mathrm{NFZ}}}$, $d_{\min,\mathrm{NFZ}}$ is the minimum safe distance between the DCU and NFZs, which is used for the DCU to adjust the speed to avoid collisions, and $d_{\mathrm{safe},\mathrm{NFZ}}$ is the safe distance threshold between the DCU and NFZs.
@@
where $d_{t,\mathbb{A}}=\|L_{\mathrm{DCU},t}-\mathbb{A}\|$ is the distance between the DCU and the boundary of $\mathbb{A}$ in the time slot $t$, $d_{\min,\mathbb{A}}$ is the minimum safe distance between the DCU and the boundary of $\mathbb{A}$, which is used for the DCU to adjust the speed to avoid flying out, and $d_{\mathrm{safe},\mathbb{A}}$ is the safe distance threshold between the DCU and the boundary of $\mathbb{A}$.
@@
where $E_{\min}(t)$ is the minimum energy required for DCU to reach the $\mathbb{LA}$ at the beginning of time slot $t$ and $E_{\min}(t)=e_{\text{fly}}(t)\cdot d_{t,\mathbb{LA}}/v_{\max}$.
$r_{9t}$ is used to encourage the DCU can complete the data collection quickly to reduce energy consumption and is as follows:
@@ 5.2 SAC Algorithm
SAC is an off-policy DRL method based on the maximum-entropy framework [35, 36]. Its off-policy property enables efficient training and easy incorporation of LLM guidance, making it suitable for continuous-control problems such as UAV trajectory planning.
SAC maximizes both the expected cumulative reward and the policy entropy:
$\pi^{*}=\arg\max_{\pi}\sum_{t=0}^{\infty}\mathbb{E}_{\mathcal{P},\pi}\!\left[\delta^{t}\big(r_{t}+\mu\,\mathcal{H}(\pi(\cdot\mid o_{t}))\big)\right],$ (20)
where $\delta$ is the discount factor, $\mu$ is the temperature parameter, and $\mathcal{H}(\pi(\cdot\mid o_{t}))=-\int\pi(a_{t}\mid o_{t})\log\pi(a_{t}\mid o_{t})\,da_{t}$ is the policy entropy. The soft Bellman equation is defined as:
$\left\{\begin{aligned} Q_{\text{soft}}(o_{t},a_{t})&=r_{t}+\delta\,\mathbb{E}_{\mathcal{P}}\!\left[V_{\text{soft}}(o_{t+1})\right],\\ V_{\text{soft}}(o_{t})&=\mathbb{E}_{a_{t}\sim\pi}\!\left[Q_{\text{soft}}(o_{t},a_{t})-\mu\log\pi(a_{t}\mid o_{t})\right].\end{aligned}\right$. (21)
The critic network minimizes the following loss:
$J_{Q}(\varphi_{i})=\mathbb{E}_{(o_{t},a_{t},r_{t},o_{t+1})\sim D}\!\left[\big(Q_{i}^{\theta}(o_{t},a_{t})-y_{t}\big)^{2}\right],$ (22)
with target value $y_{t}=r_{t}+\delta\!\left(\min_{j}Q_{j}^{\theta_{\text{tar}}}(o_{t+1},\tilde{a}_{t+1})-\mu\log\pi^{\phi}(\tilde{a}_{t+1}\mid o_{t+1})\right)$, where $D$ is the experience pool, and $\theta_{\text{tar}}$ denotes target critic parameters. The actor is optimized via reparameterization:
$J_{\pi}(\phi)=\mathbb{E}_{o_{t}\sim D,\,\varsigma\sim\mathcal{K}(0,1)}\Big[$ $\,\mu\log\pi^{\phi}(\tilde{a}_{t}^{\phi}(o_{t},\varsigma)\mid o_{t})$ (23)
$-Q^{\theta}(o_{t},\tilde{a}_{t}^{\phi}(o_{t},\varsigma))\Big],$
where $\tilde{a}_{t}^{\phi}(o_{t},\varsigma)=\tanh\!\big(\beta^{\phi}(o_{t})+\alpha^{\phi}(o_{t})\odot\varsigma\big)$, $\beta^{\phi}$ and $\alpha^{\phi}$ denote the Gaussian mean and standard deviation, $\varsigma$ is a standard normal noise, and $\odot$ is element-wise multiplication.
Through iterative updates of $\pi$, $Q$, and $V$, SAC achieves a stable, high-performance policy for continuous control under uncertainty.
@@ 5.3.1 Main Idea
Based on the POMDP formulation in Section 5.1 and the continuous nature of DCU’s velocity, we adopt SAC algorithm to optimize the DCU’s flight trajectory during the data collection. However, in the context of low-altitude economy scenarios, where strict regulatory compliance and robust obstacle-avoidance capability are critical, the use of DRL alone may be insufficient due to its reliance on trial-and-error learning and lack of prior domain knowledge. To address this issue, we propose an enhanced decision-making algorithm by integrating an LLM into the action selection module, as illustrated in Fig.  2. Specifically, when the DCU detects the presence of nearby obstacles (e.g., BZs, NFZs, RZs, or OUs), a decision condition is triggered that activates the LLM-based action selection pathway. In this mode, a Prompt Generator compiles semantic information, including environmental context, DCU status, and obstacle details, into a structured prompt, which is then input to an LLM. The LLM responds with a recommended action, leveraging its rich prior knowledge to ensure safety and regulatory compliance. An Extractor then parses the LLM’s response into actionable control parameters for the DCU. Conversely, when no immediate obstacle is detected, the system defaults to the standard SAC policy network for action selection. This hybrid structure enables the agent to benefit from both data-driven learning and knowledge-driven reasoning: the SAC component provides efficient trajectory planning in general scenarios, while the LLM supplements decision-making in critical or unfamiliar environments requiring sophisticated judgment. By combining the generalization capability of DRL with the reasoning capability and prior knowledge of an LLM, the proposed algorithm achieves superior performance in complex real-world scenarios, ensuring not only energy efficiency but also operational safety and compliance.
(a) Task definition and observation description: you are a reinforcement learning decision agent controlling a Data Collection UAV (DCU) in a 500×500 m low-altitude environment. The primary objective is to ensure safety and avoid collisions. The secondary objective is to maintain compliance—never enter NFZs, and if inside an RZ, do not exceed the local speed limit 5m/s. It can observe all GEs and their remaining data, static obstacles (buildings, NFZs, RZs), and only the nearby dynamic obstacles (OUs) within the observation radius. Decide the next continuous velocity vector $(v_{x},v_{y})$ for the DCU. The observation includes the DCU state (position, velocity), GEs (positions and remaining data), static obstacles (positions of buildings, NFZs, and RZs), and dynamic obstacles (visible OUs and positions). (b) Chain-of-thought reasoning guidelines: 1) Safety first: ensure $(v_{x},v_{y})$ will not cause collisions with buildings, NFZ borders, or visible OUs; 2) Compliance: never enter NFZs and within RZs ensure $|v|\leq\text{rz\_limit}$; 3) Data efficiency: move toward GEs, which have remaining data; 4) Energy efficiency: prefer smaller $|v|$ if equally effective; 5) Stability: maintain current direction when options are equivalent. (c) Action space and output format: the action is a continuous 2D velocity $(v_{x},v_{y})$, each within [10,10][-10,10]. The magnitude $|v|=\sqrt{v_{x}^{2}+v_{y}^{2}}$ represents the flight speed (m/s), where +x denotes east and +y denotes north. The model must output a strict JSON object as follows:
{"vx":4.2,"vy":1.6,"Confidence":0.86,"Reasoning":"Moves toward GE with remaining data, avoids NFZ/OUs/buildings, and respects RZ."}
Figure 3: CoT-based prompt generator used for UAV control. Part (a) defines the task and observation, (b) describes the reasoning principles, and (c) specifies the action format and compact JSON output.
@@ 5.3.2 Algorithm Detail
Our proposed algorithm, which is detailed in Algorithm 1, is designed to enable the DCU to effectively perform trajectory planning by learning an optimal velocity control policy that guides its trajectory to maximize data acquisition and avoid collision. The main process of our algorithm is shown in Fig. 2.
@@
After the hybrid SAC–LLM training process described above, we further detail the design of the LLM Prompt Generator, which enables the integration of LLM-based reasoning into the decision loop. When the decision condition is triggered, the prompt generator converts the UAV’s structured observation into a semantically rich text prompt that guides the LLM’s inference. As shown in Fig. 3, the generator follows a chain-of-thought (CoT) design composed of three tightly connected parts: (i) task definition and observation description, which specifies the UAV’s operating environment and observation elements; (ii) reasoning guidelines, which instruct the LLM to reason according to safety, compliance, and efficiency rules; and (iii) output format, which constrains the model to produce a strict JSON object representing a continuous velocity vector $(v_{x},v_{y})$, confidence score, and reasoning summary. This prompt structure ensures interpretability and physical validity of the LLM’s output, enabling seamless conversion into executable control actions for the UAV. Additionally, to address the potential issue of hallucination by LLM, which can lead to the failure of producing valid actions, an Extractor is employed. The extractor first checks whether any action is produced. If no action is generated, the LLM is called again, with a threshold of 10 attempts. If this threshold is exceeded, the episode is terminated. Furthermore, after receiving the action, the extractor verifies whether the magnitude of the velocity vector is less than $v_{\max}$. If the velocity exceeds $v_{\max}$, the LLM is called again, with the same 10-attempt threshold. If the limit is exceeded again, the episode is terminated. This process ensures the robustness of the decision-making loop and maintains valid and feasible UAV control actions.